# Demonstration of Efficient Radon Removal by Silver-Zeolite in a Dark Matter Detector

### v2

@@ Abstract
The regeneration process (or activation) involves flushing the trap with nitrogen for 20 minutes, while heating it to 160C using a heating tape. The temperature is monitored with a thermocouple attached to the exterior surface of the trap. After flushing, the trap is vacuumed for another 20 minutes while maintaining the temperature at 160C. Once this process is completed, we stop the heating and maintain the vacuum in the trap until it cools to room temperature. Finally, the valves are closed to isolate the trap from the circulation system. This process is used for both silver-zeolite and activated charcoal.
II S2. Analysis details
In this section, we provide details on the data selection criteria with applied quality cut and 214Po cut, and describe the methodology used to determine the expected decay rates of radon and 214Po.
II.1 Quality cut
We apply a quality cut that includes a loose pulse shape discrimination cut to remove non-physical events with characteristically short risetimes [NEWS-G:2017pxg], as well as requiring pulse amplitudes above 1200 ADU (arbitrary digital units). The latter selection removes most of the background events from atmospheric muons, as well as additional electronic noise events. Some degraded alpha events that do not deposit their full kinetic energy also fall below this cut. However, the fraction of affected alpha events remains mostly constant over time, tending to increase slightly as the detector gain decreases. A separate study quantifying the impact of this effect on signal efficiency due to this cut shows a drop of only $\sim$ 1% over time (see Fig. 3.59 in Ref. [Durnford:2024]).
Figure S1: Amplitude distribution of events for approximately 6 hours of data in phase III of campaign 1, recorded by the SPC (black markers). The distribution is modeled with an adaptive-bandwidth Gaussian kernel density estimation [silverman, wang2011bandwidth, kde_article, kde_code] with bootstrapped [efron] statistical uncertainty (red curve and $1\sigma$ shaded band). The inset shows the amplitude peak positions corresponding to the maximum energy deposition of three alpha decays versus the corresponding alpha energies (red markers with error bars), with a linear fit (blue curve and shaded band).
The remaining constant background in phase I for all campaigns (see Fig. 2 in the main text) comes predominantly from the alpha decay of 210Po (5.3 MeV [nndc_206]), originating from radioisotope depositions on the inner surface of the detector material. The SPC, made of stainless steel, does not have particular precautions to prevent 210Po contamination, but can nonetheless serve as a useful low-rate calibration source for this analysis. In phase IV, following the opening of the trap, the rate in campaigns 1 and 2 decreases significantly over several hours, eventually reaching the 210Po alpha background level observed in phase I. For campaign 3, using activated charcoal, the rate remains up to two orders of magnitude higher compared to the rate in phase I and exhibits exponential decay, indicating the continued presence of radon.
@@ II.1 Quality cut
In phases II and III, with the injection of radon, the rate increases. By applying the quality cut, we select primarily the full energy alpha decays of  [^2] Rn (5.5 MeV [nndc_218]) and its daughter isotopes, 218Po (6.0 MeV [nndc_214]) and 214Po (7.7 MeV [nndc_210]), which have clear amplitude peaks (above 12,000 ADU) in the recorded data, as shown in Fig. S1 for phase III. The nominal radon trap analysis described in the main text is performed on the data with only this quality cut applied, incorporating events of all three alpha decay isotopes; as these decays are in secular equilibrium, the proportion of events due to radon itself is approximately constant throughout each measurement campaign. Additionally, high energy beta decay events from the  [^2] Rn decay chain (214Bi in particular [nndc_214]) are present in this data selection, but are also in equilibrium with the rest of the decay chain.
@@
To verify the linear energy response of the SPC in this regime, amplitude spectra—such as the one shown in Fig. S1—are characterized with an adaptive-bandwidth Gaussian kernel density estimation [silverman, wang2011bandwidth, kde_article, kde_code], with statistical uncertainties being estimated using a bootstrapping procedure [efron]. The resulting peak amplitudes are well-fit by a linear function of the corresponding alpha decay energies, as shown in the inset of Fig. S1. A peak corresponding to 210Po might be present in this amplitude spectrum, but its rate is more than two orders of magnitude lower than the  [^2] Rn peak with which it overlaps.
II.2 214Po cut
@@
In our analysis, we develop a method to select only 214Po events. 214Po decays achieve secular equilibrium with  [^2] Rn due to its short half-life (164 $\upmu$}s [nndc_210]) and the short half-lives of the other intermediate isotopes compared to  [^2] Rn and its subsequent daughter isotopes. By selecting only 214Po events, the signal-to-background ratio is significantly improved, as there is no 210Po background in this amplitude regime.
@@
The 214Po cut applied in phases II and III is identified by the time-evolving minima between the 218Po and 214Po peaks (e.g. see Fig. S1), within approximately 6-hour data segments. This method cannot be applied directly to phase I of all campaigns and phase IV of campaigns 1 and 2, since there is no 214Po amplitude peak feature. Instead, during these phases, the 210Po amplitude peak is identified within similar 6-hour data segments, fitted with a linear function over time, and scaled to match the 214Po peak position in phases II/III, assuming energy linearity in this regime. Figure S2 illustrates the evolution of the 214Po amplitude cut across all phases for campaign 1.
Due to the large amount of radon remaining in phase IV of campaign 3 (see Fig. 2 in the main text), the 214Po amplitude cut is determined using a similar method as in phases II and III. To obtain a time-dependent 214Po amplitude cut for phase IV, the amplitude minima between the 218Po and 214Po peaks are found in approximately 6-hour data segments. The bootstrap resampling [efron] is then applied to estimate the statistical uncertainty for each point. These points are then fit with a linear function to define the 214Po amplitude cut for phase IV.
@@
Figure S2: 2D histogram of the amplitude distribution over time in campaign 1 with the trend of 214Po cut obtained across all phases. In phases II and III, the 214Po cut is defined as the amplitude minima between the 218Po and 214Po peaks in time bins of approximately 6 hours duration, indicated as red markers with statistical error bars. In phases I and IV, a linear fit of the 210Po peak position over time is scaled to match the 214Po cut at the beginning of phase II/end of phase IV, shown as red curves with shaded $1\sigma$ uncertainties.
II.3 Expected decay rates
@@
Evaluating the performance of the adsorbent requires determining the expected decay rates of radon and its daughters over time without the intervention of the trap. However, operating the SPC at event rates near 100 Hz with a 2 ms event acquisition time window introduces a significant probability of coincidental events, or pileup, and causes dead-time losses due to limitations in the data acquisition (DAQ) system. These effects reduce the observed decay rate compared to the true rate and introduce potential sources of systematic error in subsequent analyzes. It is therefore necessary to infer the true radon rate from the measured rate, accounting for pileup and dead-time losses.
@@
To achieve this, a Monte Carlo (MC) simulation of radon decay is produced for a given radon rate in addition to a given background rate, which includes the expected pileup and dead-time effects to reflect the actual conditions of our experiment. Any simulated decays occurring within 1 ms of each other (half the duration of a single event window) are treated as coincident events (resulting in a single recorded event). Simulated decays occurring within 1 to 2 ms of a preceding pulse are considered as lost due to DAQ deadtime. This approach accounts for the DAQ software centering the maximum amplitude of the triggered event trace within its acquisition time window, ensuring that no two pulses can be more than half an event window apart and still be contained in the event window. High-statistics MCs are produced for true radon rates ranging from approximately 30 Hz to 200 Hz, along with the measured 210Po background rate for each campaign described in the main text. Each MC dataset spans a duration longer than the real data from phase III, covering both phase III and phase IV. The resulting log-scale rate trends—separated into single decay event vs. pileup events—are characterized using univariate splines [scipy_spline].
The results of the decay rate MCs are incorporated into a joint model for the observed event rate over time, with and without the application of the 214Po cut. The free parameters of the model define: (i) the true radon decay rate from the discrete sets of values for which MCs are produced; (ii) the overall alpha event signal efficiency with the quality cut applied ($f_{\alpha}$); (iii) signal efficiency of observed 214Po events with the 214Po cut applied ($f_{\text{Po-214}}$); (iv) the fraction of pileup events that pass the quality or 214Po cut ($f_{\text{pileup}}$ and $f_{\text{pileup, Po-214}}$, respectively). Adding the background rates for each data selection ($\eta$ and $\eta_{\text{Po-214}}$), the expected event rates over time $\nu$, with either the quality or 214Po cut, are given as
@@
$\nu$ $(t)=f_{\alpha}\left[10^{y_{\text{Rn-222}}(t)}+10^{y_{\text{Po-218}}(t)}+10^{y_{\text{Po-214}}(t)}\right]{\color[rgb]{1,1,1}\frac{|}{|}}$
$+f_{\text{pileup}}\times 10^{y_{\text{pileup}}(t)}+\eta,{\color[rgb]{1,1,1}\frac{|}{|}}$
$\nu$ $_{\text{Po-214}}(t)=f_{\text{Po-214}}\times 10^{y_{\text{Po-214}}(t)}{\color[rgb]{1,1,1}\frac{|}{|}}$
$+f_{\text{pileup,\,Po-214}}\times 10^{y_{\text{pileup}}(t)}+\eta_{\text{Po-214}},$ (S1)
@@
A joint, binned, Poisson log likelihood function is derived for the event rate over time given this model;
$\log$ $\mathcal{L}\ =\sum_{i}\left[h_{3}(i)\ \log\nu(t_{i})-\nu(t_{i})\right]{\color[rgb]{1,1,1}\frac{|}{|}}$
$+\sum_{i}\left[h_{3}^{\text{Po-214}}(i)\ \log\nu_{\text{Po-214}}(t_{i})-\nu_{\text{Po-214}}(t_{i})\right],$ (S2)
where $h_{3}(i)$ and $h_{3}^{\text{Po-214}}(i)$ are the observed event counts in phase III in time bin $i$ (with central time $t_{i}$) with the quality cut and 214Po cut applied, respectively. This model is fit to the data using a Markov Chain Monte Carlo (MCMC). Specifically, the python package emcee is used with the “stretch-move” Metropolis-Hastings algorithm [GW10, emcee], propagating the 100 random walkers for $2\times 10^{5}$ steps after an initial burn-in phase of $4\times 10^{4}$ steps. This exceeds the recommendation that the number be at least $50\times$ the estimated autocorrelation time [emcee]. The resulting fits (with statistical uncertainty in the form of random MCMC samples) are shown for each campaign in Fig. 2 in the main text.
III S3. Rates in phases I and IV
For the three campaigns, we compare the rates obtained in phases I and IV, with either the quality or 214Po cut. The rate in phase I for all campaigns is approximately constant and is determined by the average count of all time bins in this phase. Similarly, the constant rate in phase IV of campaigns 1 and 2 is calculated starting from the reference time (vertical dashed line in Fig. 2 of the main text). However, in campaign 3, the rate decreases exponentially. To facilitate comparison with campaigns 1 and 2, we also determine the rate in phase IV at the reference time for all campaigns. Table S1 summarizes the rate in phases I and IV with the quality and 214Po cuts, as well as the rate in phase IV for all campaigns, calculated at the reference time.
Phase I rate (Hz) Phase IV rate (Hz)

### v3

@@ 2.1 Experimental setup
Before each measurement, the trap undergoes thermal regeneration at 160C within the closed-loop system. This regeneration, also known as activation, removes any adsorbed gases and contaminants. It consists of flushing the trap with ultra-high-purity nitrogen gas for 20 minutes via the gas inlet to the exhaust, while heating it to 160C using a heating tape. The temperature is monitored with a thermocouple attached to the exterior surface of the trap. The nitrogen flow rate is controlled by adjusting the nitrogen supply pressure to 15–20 PSI using the regulator on the gas cylinder. After flushing, the trap is evacuated for another 20 minutes while maintaining the temperature at 160C. Once this process is completed, we stop the heating and maintain the vacuum in the trap until it cools to room temperature. Finally, the valves are closed to isolate the trap from the circulation system. This process is used for both silver-zeolite and activated charcoal. To further minimize contaminants in the gas loop, the entire system is evacuated to approximately $10^{-7}$ Torr, except the circulation pump, which cannot achieve a vacuum lower than $10^{-3}$ Torr. Finally, the gas mixture is injected, filling the system to a pressure of 500 mbar.
During the tests, the closed‑loop system remains sealed and leak‑tight, and the adsorbents (silver‑zeolite and activated charcoal) are never exposed to ambient air. The nitrogen used to regenerate the trap is ultra‑high‑purity grade 5 gas (N2 > 99.999%), with impurity levels of H2O < 3 ppm, O2 < 3 ppm and hydrocarbon < 0.1 ppm (Linde, Part #: NI 5.0UH-T). The argon/methane mixture is a custom pre‑mix supplied by Linde; it is a non‑certified blend prepared from grade 5 argon and methane, with typically H2O < 10 ppm.
2.2 Experimental procedure
We conduct three experimental campaigns to test the efficacy of the radon trap at room temperature and demonstrate reproducibility: two with silver-zeolite (campaigns 1 and 2) and one using activated charcoal (campaign 3). To ensure consistency across the campaigns, we organize the data collection into four phases:
  • $\bullet$
    Background (phase I): initially, we perform at least a 24-hour background run to determine the minimal sensitivity of the detector;
  • $\bullet$
    Radon diffusion (phase II): radon is injected by diffusion exclusively into the detector for a day until reaching a raw event rate of $\sim$ 75 Hz. Once this rate is achieved, we close the radon source and isolate the detector. Continuing diffusion of radon from the gas system eventually raises the rate to $\gtrsim$ 100 Hz. During this period, the rest of the system remains closed and there is no circulation;
  • $\bullet$
    Radon decay (phase III): we monitor the radon decay rate for approximately two days, while the rest of the system remains closed and there is no circulation;
  • $\bullet$
    Trap open (phase IV): the closed-loop system, including the trap, is opened to the detector and the pump is used to circulate the gas, as shown in Fig. 1. Radon is removed at this step.
The four phases for the three campaigns are shown in Fig. 2. The same procedure was used and the same trap was filled with 10 g of silver zeolite or activated charcoal adsorbent for each campaign.
Figure 2: Measured rate with either quality cut only or with also the 214Po cut applied during the three radon campaigns, with the trap filled with either silver-zeolite or activated charcoal, at room temperature. The blue distribution corresponds to the background run (phase I), orange is the radon diffusion in the SPC (phase II), green is the radon decay (phase III), and red is the events recorded when the radon trap is open (phase IV). The green curves are the fitted/extrapolated decay rates for each campaign, depicting many random Markov Chain Monte Carlo (MCMC) samples. The light and dark shades make the distinction between quality cut only and with also the 214Po cut applied, respectively. The vertical dashed line defines the reference time where the expected rate with quality cut is equal to 76.5 Hz. Note that there are some visible gaps in the data (campaigns 2 and 3) due to brief pauses in data collection.
3 Data analysis methods
To interpret the data and assess the experimental results, data selection cuts are applied, and the rate of radon events over time is determined. The performance of the trap is evaluated by comparing the expected rate from phase III without the intervention of the trap and the rate observed in phase IV.
3.1 Quality cut
@@
We apply a quality cut that includes a loose pulse shape discrimination cut to remove non-physical events with characteristically short risetimes NEWS-G:2017pxg , as well as requiring pulse amplitudes above 1200 ADU (arbitrary digital units). The latter removes most of the background events from atmospheric muons, as well as additional electronic noise events, selecting primarily the full energy alpha decays of  [^2] Rn (5.5 MeV nndc_218 ) and its daughter isotopes, 218Po (6.0 MeV nndc_214 ) and 214Po (7.7 MeV nndc_210 ), which have clear amplitude peaks above 12,000 ADU (see Fig. 3). The rate over time in all three campaigns with this cut applied is shown in Fig. 2. As these three alpha-decay processes quickly reach secular equilibrium, the proportion of events due to radon itself is approximately constant throughout each measurement campaign. Additionally, high energy beta decay events from the  [^2] Rn decay chain (214Bi in particular nndc_214 ) are present in this data selection, but are also in equilibrium with the rest of the decay chain.
Some degraded alpha events that do not deposit their full kinetic energy also fall below this quality cut. However, the fraction of affected alpha events remains approximately constant—and small—over time, increasing slightly as the detector gain decreases. A separate study quantifying the impact of this effect on signal efficiency due to this cut shows a drop of only $\sim$ 1% over time (see Fig. 3.59 in Ref. Durnford:2024 ).
@@
Figure 3: Amplitude distribution of events for approximately 6 hours of data in phase II of campaign 1, recorded by the SPC (black markers). The distribution is modeled with an adaptive-bandwidth Gaussian kernel density estimation silverman ; wang2011bandwidth ; kde_article ; kde_code with bootstrap-estimated efron statistical uncertainty (red curve and $1\sigma$ shaded band). The inset shows the amplitude peak positions corresponding to the maximum energy deposition of three alpha decays versus the corresponding alpha energies (red markers with error bars), with a linear fit (blue curve and shaded band). A peak corresponding to 210Po might be present, but its rate is more than two orders of magnitude lower than the  [^2] Rn peak with which it overlaps.
The constant background remaining in phase I for all campaigns (see Fig. 2) comes predominantly from the alpha decay of 210Po (5.3 MeV nndc_206 ), originating from radioisotope depositions on the inner surface of the detector material. The SPC, made of stainless steel, does not have particular precautions to prevent 210Po contamination, but can nonetheless serve as a useful low-rate calibration source for this analysis. In phase IV, following the opening of the trap, the rate in campaigns 1 and 2 decreases significantly over several hours, eventually reaching the 210Po alpha background level observed in phase I. For campaign 3, using activated charcoal, the rate remains up to two orders of magnitude higher compared to the rate in phase I and exhibits exponential decay, indicating the continued presence of radon.
3.2 214Po cut
@@
To achieve a lower background rate than what remains after applying the quality cut, an additional stricter selection is applied to select only 214Po events. 214Po decays achieve secular equilibrium with  [^2] Rn due to its short half-life (164 $\upmu$}s nndc_210 ) and the short half-lives of the other intermediate isotopes compared to  [^2] Rn and its subsequent daughter isotopes. As shown in Fig. 2, the selection of only 214Po events significantly improves our signal-to-background ratio and allows us to assess our results with near-zero background in later analyses, as there is no 210Po background in this amplitude regime.
@@
When radon is injected in phases II and III, the amplitude distribution exhibits different peaks corresponding to  [^2] Rn (5.5 MeV nndc_218 ) and the daughter isotopes 218Po (6.0 MeV nndc_214 ) and 214Po (7.7 MeV nndc_210 ), proportional to the energy of each alpha emitter (see Fig. 3). The 214Po cut applied is identified by the time-evolving minima between the 218Po and 214Po peaks, within approximately 6-hour data segments. This method cannot be applied directly to phase I of all campaigns and phase IV of campaigns 1 and 2, since there is no 214Po amplitude peak feature. Instead, during these phases, the 210Po amplitude peak is identified within similar 6-hour data segments, fit with a linear function over time, and scaled to match the amplitude corresponding to the 214Po cut defined in phases II and III, assuming energy linearity in this regime. An example of the evolution of the 214Po amplitude cut across all phases for campaign 1 is shown in Fig. 4.
@@
Figure 4: 2D histogram of the amplitude distribution over time in campaign 1 with the trend of 214Po cut obtained across all phases. In phases II and III, the 214Po cut is defined as the amplitude minima between the 218Po and 214Po peaks in time bins of approximately 6 hours duration, indicated as red markers with statistical error bars. In phases I and IV, a linear fit of the 210Po peak position over time is scaled to match the 214Po cut at the beginning of phase II/end of phase III, shown as red curves with shaded $1\sigma$ uncertainties.
To verify the linear energy response of the SPC in this regime assumed to perform the scaling described above, amplitude spectra—such as the one shown in Fig. 3—are characterized with an adaptive-bandwidth Gaussian kernel density estimation silverman ; wang2011bandwidth ; kde_article ; kde_code , with statistical uncertainties being estimated using a bootstrapping procedure efron . The resulting peak amplitudes are well-fitted by a linear function of the corresponding alpha decay energies, as shown in the inset of Fig. 3.
Due to the large amount of radon remaining in phase IV of campaign 3 (see Fig. 2), the 214Po amplitude cut is determined using a similar method as in phases II and III. To obtain a time-dependent 214Po amplitude cut for phase IV, the amplitude minima between the 218Po and 214Po peaks are found in approximately 6-hour data segments. The bootstrap resampling efron is then applied to estimate the statistical uncertainty for each point. These points are then fit with a linear function to define the 214Po amplitude cut for phase IV.
3.3 Expected decay rates
@@
Evaluating the performance of the adsorbent requires determining the expected decay rates of radon and its daughters over time without the intervention of the trap. The expected decay rates with either just the quality cut or also with the 214Po cut applied are determined using the phase III data of each campaign. However, operating the SPC at event rates near 100 Hz with a 2 ms event acquisition time window introduces a significant probability of coincidental events, or pileup, and causes dead-time losses due to limitations in the data acquisition (DAQ) system Durnford:2024 . These effects reduce the observed decay rate compared to the true rate and introduce potential sources of systematic error in subsequent analyzes. It is therefore necessary to infer the true radon rate from the measured rate, accounting for pileup and dead-time losses.
@@
To achieve this, a Monte Carlo (MC) simulation of radon decay is produced for a given radon rate in addition to a given background rate, which includes the expected pileup and dead-time effects to reflect the actual conditions of our experiment. Any simulated decays occurring within 1 ms of each other (half the duration of a single event window) are classified as coincident events (resulting in a single recorded event). Simulated decays occurring within 1 to 2 ms of a preceding pulse are considered lost due to DAQ deadtime. This treatment accounts for the DAQ software centering the maximum amplitude of the triggered event trace within its acquisition time window, ensuring that no two pulses can be more than half an event window apart and still be contained in the event window. High-statistics MCs are produced for true radon rates ranging from approximately 30 Hz to 200 Hz, along with the measured 210Po background rate (phase I) for each campaign. Each MC dataset spans a duration longer than the real data from phase III, covering the time-spans of both phase III and phase IV. The resulting log-scale rate trends—separated into single decay event vs. pileup events—are characterized using univariate splines scipy_spline .
The results of the decay rate MCs are incorporated into a joint model for the observed event rate over time, with and without the application of the 214Po cut. The free parameters of the model define:
  • (1)
    the true radon decay rate at the start of each phase III fit, from the discrete sets of values for which MCs are produced;
  • (2)
    the overall alpha event signal efficiency with the quality cut applied ($f_{\alpha}$);
  • (3)
    signal efficiency of observed 214Po events with the 214Po cut applied ($f_{\text{Po-214}}$);
  • (4)
    the fraction of pileup events that pass the quality or 214Po cut ($f_{\text{pileup}}$ and $f_{\text{pileup, Po-214}}$, respectively).
Adding the background rates for each data selection ($\eta$ and $\eta_{\text{Po-214}}$), the expected event rates over time $\nu$, with and without the 214Po cut, are given as
@@
$\nu$ $(t)=f_{\alpha}\left[10^{y_{\text{Rn-222}}(t)}+10^{y_{\text{Po-218}}(t)}+10^{y_{\text{Po-214}}(t)}\right]+f_{\text{pileup}}\times 10^{y_{\text{pileup}}(t)}+\eta,{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\frac{|}{|}}$
$\nu$ $_{\text{Po-214}}(t)=f_{\text{Po-214}}\times 10^{y_{\text{Po-214}}(t)}+f_{\text{pileup,\,Po-214}}\times 10^{y_{\text{pileup}}(t)}+\eta_{\text{Po-214}},$ (1)
@@
A joint, binned, Poisson log-likelihood function is derived for the event rate over time given this model;
$\log$ $\mathcal{L}\ =\sum_{i}\left[h_{3}(i)\ \log\nu(t_{i})-\nu(t_{i})\right]+\sum_{i}\left[h_{3}^{\text{Po-214}}(i)\ \log\nu_{\text{Po-214}}(t_{i})-\nu_{\text{Po-214}}(t_{i})\right],$ (2)
where $h_{3}(i)$ and $h_{3}^{\text{Po-214}}(i)$ are the observed event counts in phase III in time bin $i$ (with central time $t_{i}$) with and without the 214Po cut applied, respectively. This model is fit to the data using a Markov Chain Monte Carlo (MCMC). Specifically, the python package emcee is used with the “stretch-move” Metropolis-Hastings algorithm GW10 ; emcee , propagating the 100 random walkers for $2\times 10^{5}$ steps after an initial burn-in phase of $4\times 10^{4}$ steps. This exceeds the recommendation that the number be at least $50\times$ the estimated autocorrelation time emcee . The resulting fit of the observed event rates with each cut applied is shown in Fig. 2, including extrapolations to phase IV for each campaign to show the expected decay rates without the intervention of the radon trap.
4 Results and discussion
4.1 Rate comparison
Figure 2 clearly demonstrates that campaigns 1 and 2, which utilize the silver-zeolite trap, outperform campaign 3 using activated charcoal in terms of radon and 214Po event reduction. Indeed, the rates in phase IV drop significantly before stabilizing. In contrast, the drop in rate is much smaller in phase IV of campaign 3, with an exponential decay indicating the presence of radon remaining in the detector.
As an initial assessment of the trap performance, we compare the rates obtained in phases I and IV with and without the 214Po cut throughout the three campaigns. As can be seen in Fig. 2 the rate in phase I for all campaigns is approximately constant, and is fitted with a constant function. To compare this rate with the rate in phase IV after the trap is opened, we define the following reference time. We extrapolate the expected rate from phase III (without trap) with quality cut to later times and determine the reference time where the extrapolation reaches 76.5 Hz (dashed vertical lines in Fig. 2). Then, the constant rate in phase IV of campaigns 1 and 2 is determined by the average count of all time bins starting from the reference time. However, in campaign 3, the rate decreases exponentially and cannot be fitted with a constant function. To facilitate comparison with campaigns 1 and 2, we also determine the rate in phase IV at the reference time for all campaigns. Table 1 summarizes the rate in phases I and IV with the quality and 214Po cuts, as well as the rate in phase IV for all campaigns, calculated at the reference time. From this, we calculate the rate ratio between phase IV and phase I with either the quality or 214Po cut, as shown in Fig. 5.
Table 1: Rate in phase I (background) and phase IV (trap open) with either the quality or 214Po cut for all campaigns. The rate in phase IV obtained from the constant fit and at the reference time are also included for all campaigns.
Rate (Hz) Phase I Phase IV a Phase IV b